OpenStudy (anonymous):

What is 5 pi over 4 radians in degrees? 45° 135° 225° 315°

OpenStudy (nipunmalhotra93):

7pi/4=2pi-pi/4.... so you can evaluate all three at -pi/4 as f(2kpi+x)=f(x) where f is a trig function

OpenStudy (anonymous):

so, how is 7pi/4=2pi-pi/4 2pi - pi is pi, right?

OpenStudy (nipunmalhotra93):

\[\frac{ 7 \Pi }{ 4 }=2\Pi-\frac{ \Pi }{ 4 }\]

OpenStudy (nipunmalhotra93):

^that's what I meant.

OpenStudy (anonymous):

actually.... nevermind.. i figured it out O.o thanks for your help though :D

OpenStudy (nipunmalhotra93):

you're welcome.